Session: May 2001 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Mon Jun 4 11:54:49 METDST 2001) Success

Torun (K.Borkowski at Mon Jun 4 12:46:04 METDST 2001) Minor failures
BBC5 and BBC8 faultering (very low levels; occasionally correct) throughout the experiment. No other known problems.

Noto (P. Leto at Mon Jun 4 14:08:17 METDST 2001) Minor failures
BBC8 has variable LSB levels, at 8:20 UT of DOY 155 this problem has been solved

Jodrell_Bank (Alastair Gunn at Mon Jun 4 18:33:54 METDST 2001) Success
No known problems.

Cambridge (Alastair Gunn at Mon Jun 4 18:34:33 METDST 2001) Minor failures
Problems with final tape starting at 04:04 UT. Started recording at 05:06 UT but had frequent loss of vacuum (see log for details). Recorder servo checked out and okay. Probably a bad tape.

Onsala (John Conway at Tue Jun 5 11:09:03 METDST 2001) Success
No known problems

Westerbork (Tony Foley at Tue Jun 5 16:04:34 METDST 2001) Minor failures
Phasecal not working. Otherwise OK

Effelsberg (A. Kraus at Wed Jun 6 08:37:12 METDST 2001) Success